# Unique Security and Privacy Threats of Large Language Model: A Comprehensive Survey

### v1

@@ Abstract.
With the rapid development of artificial intelligence, large language models (LLMs) have made remarkable progress in natural language processing. These models are trained on large amounts of data to demonstrate powerful language understanding and generation capabilities for various applications, from machine translation and chatbots to agents. However, LLMs have exposed a variety of privacy and security issues during their life cycle, which have become the focus of academic and industrial attention. Moreover, these risks LLMs face are pretty different from previous traditional language models. Since current surveys lack a clear taxonomy of unique threat models based on diverse scenarios, we highlight unique privacy and security issues based on five scenarios: pre-training, fine-tuning, RAG system, deploying, and LLM-based agent. Concerning the characteristics of each risk, this survey provides potential threats and countermeasures. The research on attack and defense situations LLMs face can provide feasible research directions, making more areas reap LLMs’ benefits.
@@
copyright: acmcopyrightjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: POMACSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
With the rapid development of artificial intelligence (AI) technology, researchers have progressively expanded the scale of training data and model architectures (Zhao et al., 2023c). Learned with massive amounts of data, extremely large-scale models demonstrate impressive language understanding and generation capabilities (Naveed et al., 2023), marking a significant breakthrough in natural language processing (NLP). Referred to as Large Language Models (LLMs), those models provide strong support for machine translation, text summary, automatic coding, and other NLP tasks. However, the in-depth application of LLMs across various industries, such as chatbots (Chowdhury et al., 2023) and medical diagnosis (Richens et al., 2020), expose their life cycle to various privacy and security issues. More importantly, LLMs face unique privacy and security risks (Cui et al., 2024) that never posed in traditional language models, demanding higher requirements for privacy protection and security defenses.
@@ 1.1. Motivation
Compared to traditional single-function language models, LLMs demonstrate remarkable understanding abilities, deploying across various applications such as logical reasoning and code generation. Recently, an increasing number of companies are launching universal or domain-specific LLMs, such as ChatGPT (Chowdhury et al., 2023) and LLaMA (Touvron et al., 2023), offering users with versatile and intelligent services. However, due to LLMs’ unique power and structures, throughout their life cycle, LLMs meet with unique security and privacy threats from society compared with previous single-function models or small-scale models (Yao et al., 2024a). Existing surveys describe various risks and countermeasures by method type, and there is a lack of exploration of those unique threats. We divide LLMs’ life cycle into five scenarios and discuss the unique privacy and security risks (Das et al., 2024) at each scenario.
@@
Unique privacy risks. When learning language knowledge from the training data, LLMs would memorize the data (Carlini et al., 2022). It allows adversaries to steal private information. For example, Carlini et al. (Carlini et al., 2021) found that prompts with specific prefixes could make GPT-2 generate content containing personal information, such as email addresses and phone numbers. When running inference, the unlimited use of LLMs would provide adversaries with opportunities to steal model-related information (Zhu et al., 2022) and functionalities (Zhou et al., 2024). In summary, throughout the life cycle of LLMs, adversaries can steal or infer various sensitive information, thus threatening specific individuals and institutions.
@@
Unique security risks. Since the training data contains malicious, illegal, hallucinatory, and biased texts, LLMs inevitably learn negative language knowledge. Moreover, malicious third parties responsible for developing LLMs in outsourcing scenarios can affect these models’ integrity and utility by poisoning attacks (Zhou et al., 2022) and backdoor attacks (Wang et al., 2023a). For example, an attacker could implant a backdoor in an LLM-based automated customer service system, causing the system to respond automatically with a predetermined fraudulent link when asked specific questions. When running inference, the unlimited use of LLMs allows adversaries to obtain targeted responses (Liu et al., 2023b), such as fake news, phishing sites, and illegal content. In summary, adversaries can exploit internal and external vulnerabilities in LLMs throughout their life cycle to implement various security attacks (Yao et al., 2024a).
@@
These unique privacy and security risks would pose more severe attacks to our society, such as reducing the credibility of LLMs and hindering their application or popularity. At the same time, these risks threaten the safety of LLM owners and users, violating existing laws, such as the General Data Protection Regulation (GDPR). In this context, systematic research on LLMs’ unique privacy and security issues is lacking. It prompts us to analyze, categorize and summarize the existing research to complete a comprehensive survey in this field. This research can help the technical community to develop safe and reliable LLM-based applications, making more areas reap LLMs’ benefits.
@@ 1.2. Comparison with existing surveys
Research on LLMs’ privacy and security is rapidly developing while existing surveys need a comprehensive taxonomy and summary. As shown in Table 1, we compared our survey with the existing surveys proposed since June 2024. The main differences lie in four aspects.
@@
    Threat scenarios. We explicitly divided the life cycle of LLMs into five threat scenarios, which most surveys overlooked. Each scenario corresponds to multiple threat models, such as a pre-training scenario with malicious contributors, developers and users, like Figure 2. For each threat model, adversaries can compromise the LLMs’ privacy and security via various attack methods.
@@
    Taxonomy. We categorized the risks LLMs face based on their life cycle. Other surveys lacked such fine-grained taxonomy, making distinguishing the characteristics of various risks hard.
@@
    Unique privacy/security. Concerning privacy/security risks, we focused on the unique risks to LLMs. Meanwhile, we also explored the common risks of all language models. In response to these risks, we summarized potential countermeasures. However, most of the surveys lacked depth and comprehensiveness in this analysis, simply listing attacks and defense methods.
@@
    Other unique privacy/security scenarios. We incorporated LLMs with three scenarios: federated learning, machine unlearning and watermarking. Since these scenarios involve privacy/security issues, we explored them in Section 9 and gave a systematic study that most surveys overlooked.
@@ 1.3. Contributions of this survey
LLMs have a strong emergent ability, which is applied to many industries. However, many vulnerabilities in their life cycle pose privacy and security risks. It seriously hinders the application and extension of LLMs. Hence, we aim to analyze, categorize, and summarize privacy and security issues. Specifically, we propose a novel taxonomy for these risks, which offers a clear and comprehensive analysis of their goals, causes, and implementation methods. We hope this survey can provide researchers with feasible research directions and challenges. The main contributions of this survey are as follows.
@@
    Taking the LLMs’ life cycle as a clue, we considered risks and countermeasures in five different scenarios from those of traditional models. These scenarios include pre-training, fine-tuning, RAG system, deploying and LLM-based agent.
@@
    For each scenario, we highlighted the differences in privacy and security risks between LLMs and traditional models. Specifically, we described unique risks to LLMs and common risks to all models. Given a risk, we detailed its attack goal and capacity and combed through the related research using the attack method. Meanwhile, we analyzed the mapping between risks and countermeasures, thus listing potential defense methods.
@@
  • $\bullet$
    We provided critical and comprehensive explorations on LLMs’ privacy and security issues and pointed out feasible research directions.
@@ 2.1. Definition of LLM
LLMs represent a revolution in the field of NLP. To improve the efficiency of text processing, researchers proposed pre-trained language models based on transformers. Google released the BERT model that uses bidirectional transformers, solving downstream tasks in the ‘pre-train + fine-tune’ paradigm. Subsequently, they expanded the scale of pre-trained models, more than billions of parameters (e.g., GPT-3), and introduced novel techniques like zero-shot or few-shot learning. Compared to small-scale pre-trained models, large-scale pre-trained models showcase remarkable emergent abilities not found in these regular-scale models, capable of handling unseen tasks through in-context learning (Xu et al., 2024a) (i.e., without retraining) and instruction tuning (Shu et al., 2023) (i.e., lightweight fine-tuning).
Recent studies summarize four key characteristics that LLMs should have (Zhao et al., 2023c; Xu et al., 2024b). First, an LLM could understand various texts, like training data and natural language instructions. Second, it should be able to solve unseen NLP tasks without updating the model parameters. Third, it should generate high-quality texts that align with humans when providing designed prompts. Fourth, an LLM should showcase contextual awareness considering specific factors such as domain expertise. Moreover, Wei et al. (Wei et al., 2022) found that language models with more than 1 billion parameters have jump performance on multiple NLP tasks. Therefore, we believe that LLMs must have more than a billion parameters. Numerous institutions have developed their LLMs with these characteristics.
@@ 2.2. Traditional privacy and security risks
Recent research on privacy and security risks in artificial intelligence focuses on small-scale models.
For privacy risks, small-scale models’ life cycle contains confidential information like raw data and model details. It could lead to severe economic losses if this information is leaked (Ye et al., 2024). Raw data exposes PII, such as facial images. Reconstruction attacks (Morris et al., 2023b) and model inversion attacks (Ye et al., 2024) can steal raw data using gradients or logits. In addition, member and attribute information is sensitive. Taking medical tasks as an example, adversaries can adopt membership inference attacks (Ye et al., 2022) to judge if an input belongs to the training set, revealing some users’ health conditions. Model details have significant commercial value and are subject to model extraction attacks (Li et al., 2023e), targeting black-box victim models to obtain substitute counterparts or partial model information (e.g., network architecture and optimization algorithm) via multiple queries. Then, adversaries who know partial model details could launch stronger privacy and security attacks.
For security risks, small-scale models face poisoning attacks (Wan et al., 2023) that compromise model utility by modifying the training data. A backdoor attack is a variant of poisoning attacks (Wang et al., 2023a; Ma et al., 2023). It can inject hidden backdoors into the victim model by manipulating training data or model parameters, thus controlling the returned outputs. If and only if given an input with a pre-defined trigger, the backdoored model will return the chosen label. During inference, adversarial example attacks (Guo et al., 2021) craft adversarial inputs by adding imperceptible perturbation, making them cause incorrect predictions. In summary, these security attacks can compromise model utility and integrity and severely threaten public safety in practical applications. Taking sentiment analysis tasks on social media platforms as an example, all attacks could lead to racially discriminatory texts being classified as benign, which results in the widespread dissemination of harmful information.
The life cycle of LLMs shares similarities with and differs from small-scale models. From traditional privacy and security risks, we outline various potential risks LLMs face in different scenarios, as shown in Figure 1. Each scenario has a unique data type and implementation process, showing different risks and their countermeasures. We will start with each scenario and its threat models.
Figure 1. The pipeline of the survey. For each threat scenario, the first column is the data type used, and the second column is the process applied, where the text boxes are unique data types and processes to LLMs. The fourth column and the fifth column are extended risks and countermeasures, respectively. Where the underlined texts are the unique risks for LLMs, and other texts are the common risks of all language models.
@@ 3. Threat Scenarios for LLMs
Although many institutions have disclosed the implementation methods of their LLMs, some details are still unknown. We collect literature and divide the life cycle of an LLM into five scenarios at a finer granularity rather than only training and inference phases. Figure 1 gives these unique threat scenarios: pre-training LLMs, fine-tuning LLMs, retrieval augmented generation (RAG) system, deploying LLMs and deploying LLM-based agents. Then, we list extended risks for each scenario and use underlined texts to highlight unique risks for LLMs.
@@ 3.1. Pre-training LLMs
In this scenario, model developers collect a large corpus as a pre-trained data set containing books (Zhu et al., 2015), web pages (e.g., Wikipedia), conversational texts (e.g., Reddit), and code (e.g., Stack Exchange). Then, they use large-scale, Transformer-based networks and advanced training algorithms, enabling the models to learn rich language knowledge from vast amounts of unlabeled texts. After obtaining the pre-trained LLM, developers upload it to open-source community platforms to either gain profits or contribute to the community’s development, as shown in Figure 2. In this case, we consider three malicious entities: contributors, developers, and users.
@@
    Malicious contributors. Unlike small-scale language models, the corpora involved in pre-training LLMs are so large that developers can not audit all the data, resulting in the inevitable existence of negative texts(e.g., toxic data and privacy data). These negative texts directly affect the safety of LLMs. For example, an LLM can learn steps to make a bomb from illegal data and feed the details back to the user. In this survey, we focus on the privacy and security risks posed by toxic data and privacy data without discussing the issues of fairness and hallucination.
@@
    Malicious developers. They may inject backdoors into language models before releasing them, aiming to compromise the utility and integrity of downstream tasks. If victim users download and deploy a compromised model, the attacker who knows the trigger can easily activate the hidden backdoor, thus manipulating the compromised model.
@@
    Malicious users. After downloading public models, they gain access to their information except for the training data, effectively becoming white-box attackers. Thus, these users can perform inference and data extraction attacks in a white-box setting.
@@ 3.2. Fine-tuning LLMs
In this scenario, users would customize LLMs for specific NLP tasks. They download pre-trained LLMs from open-source platforms and fine-tune them on the customized dataset. There are three fine-tuning methods: supervised learning, instruction and alignment tuning. The first manner is the commonly used training algorithm. For the second manner, the instruction is in natural language format containing a task description, an optional demonstration, and an input-output pair (Wang et al., 2023b). Through a sequence-to-sequence loss, instruction-tuning can help LLMs understand and generalize to unseen tasks. The third manner aligns LLMs’ outputs with human preferences, like usefulness, honesty and harmlessness. To meet these goals, Ziegler et al. (Ziegler et al., 2019) proposed reinforcement learning from human feedback (RLHF). Figure 5 shows two malicious entities: contributors and third parties.
@@
    Malicious contributors. In general, users need to collect specific samples that are used to fine-tune downstream tasks. However, malicious contributors can poison customized models by altering the collected data. In this scenario, the adversary can only modify a fraction of the contributed data without access to the fine-tuning.
@@
    Malicious third-parties. When outsourcing customized LLMs, users share their local data with third-party trainers with computational resources and expertise. However, malicious trainers could poison these customized LLMs to manipulate their responses before delivering them to users. For example, in a Question-Answer task, the adversary could manipulate the customized LLM to return misleading responses (e.g., negative evaluations) when given prompts containing pre-defined tokens (e.g., celebrity names). Compared to small-scale language models, fine-tuning LLMs involves two unique techniques: instruction-tuning and alignment-tuning. Therefore, malicious trainers pose two unique risks to LLMs: poisoning instruction-tuning and RLHF. In addition, we also consider a security risk common to all language models, such as poisoning supervised learning.

### v2

@@ Abstract.
With the rapid development of artificial intelligence, large language models (LLMs) have made remarkable advancements in natural language processing. These models are trained on vast datasets to exhibit powerful language understanding and generation capabilities across various applications, including machine translation, chatbots, and agents. However, LLMs have revealed a variety of privacy and security issues throughout their life cycle, drawing significant academic and industrial attention. Moreover, the risks faced by LLMs differ significantly from those encountered by traditional language models. Given that current surveys lack a clear taxonomy of unique threat models across diverse scenarios, we emphasize the unique privacy and security threats associated with five specific scenarios: pre-training, fine-tuning, retrieval-augmented generation systems, deployment, and LLM-based agents. Addressing the characteristics of each risk, this survey outlines potential threats and countermeasures. Research on attack and defense situations can offer feasible research directions, enabling more areas to benefit from LLMs.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Security and privacy Human and societal aspects of security and privacy
@@ 1. Introduction
With the rapid development of artificial intelligence (AI) technology, researchers have progressively expanded the scale of training data and model architectures (Zhao et al., 2023c). Trained with massive amounts of data, extremely large-scale models demonstrate impressive language understanding and generation capabilities (Naveed et al., 2023), marking a significant breakthrough in natural language processing (NLP). Referred to as Large Language Models (LLMs), these models provide robust support for machine translation, text summarization, automatic coding, and other NLP tasks. However, the in-depth application of LLMs across various industries, such as chatbots (Chowdhury et al., 2023) and medical diagnosis (Richens et al., 2020), exposes their life cycle to numerous privacy and security threats. More importantly, LLMs face unique privacy and security risks (Cui et al., 2024) not present in traditional language models, necessitating higher standards for privacy protection and security defenses.
@@ 1.1. Motivation
Compared to traditional single-function language models, LLMs demonstrate remarkable comprehension abilities and are deployed across various applications, such as logical reasoning and code generation. Recently, an increasing number of companies have been launching universal or domain-specific LLMs, such as ChatGPT (Chowdhury et al., 2023) and LLaMA (Touvron et al., 2023), offering users versatile and intelligent services. However, due to LLMs’ unique capabilities and structures, they encounter unique security and privacy threats throughout their life cycle compared to previous single-function or small-scale models (Yao et al., 2024a). Existing surveys describe various risks and countermeasures by method type, but there is a lack of exploration into these unique threats. We divide the life cycle of LLMs into five scenarios and discuss the unique privacy and security risks (Das et al., 2024) inherent to each scenario.
@@
Unique privacy risks. When learning language knowledge from training data, LLMs tend to memorize this data (Carlini et al., 2022). This tendency allows adversaries to extract private information. For example, Carlini et al. (Carlini et al., 2021) found that prompts with specific prefixes could cause GPT-2 to generate content containing personal information, such as email addresses and phone numbers. When running inference, unrestricted use of LLMs provides adversaries with opportunities to extract model-related information (Zhu et al., 2022) and functionalities (Zhou et al., 2024). In summary, throughout the life cycle of LLMs, adversaries can extract or infer various sensitive information, thereby threatening specific individuals and institutions.
@@
Unique security risks. Since the training data may contain malicious, illegal, hallucinatory, and biased texts, LLMs inevitably acquire negative language knowledge. Moreover, malicious third parties involved in developing LLMs in outsourcing scenarios can compromise these models’ integrity and utility through poisoning attacks (Zhou et al., 2022) and backdoor attacks (Wang et al., 2023a). For example, an attacker could implant a backdoor in an LLM-based automated customer service system, causing it to respond with a predetermined fraudulent link when asked specific questions. When running inference, unrestricted use of LLMs allows adversaries to obtain targeted responses (Liu et al., 2023b), such as fake news, phishing sites, and illegal content. In summary, adversaries can exploit internal and external vulnerabilities in LLMs throughout their life cycle to execute various security attacks (Yao et al., 2024a).
@@
These unique privacy and security risks pose severe threats to society, such as reducing the credibility of LLMs and hindering their popularity. Additionally, these risks threaten the safety of LLM owners and users, violating existing laws such as the General Data Protection Regulation (GDPR). In this context, there is a lack of systematic research on the unique privacy and security threats of LLMs. This prompts us to analyze, categorize, and summarize the existing research to complete a comprehensive survey in this field. This research can help the technical community develop safe and reliable LLM-based applications, enabling more areas to benefit from LLMs.
@@ 1.2. Comparison with existing surveys
Research on LLMs’ privacy and security is rapidly developing, but existing surveys lack a comprehensive taxonomy and summary. As shown in Table 1, we compared our survey with existing surveys published since June 2024. The main differences lie in four key aspects.
@@
    Threat scenarios. We explicitly divided the life cycle of LLMs into five threat scenarios, which most surveys overlooked. Each scenario corresponds to multiple threat models, such as a pre-training scenario involving malicious contributors, developers, and users, as shown in Figure 2. For each threat model, adversaries can compromise the LLMs’ safety through various attack methods.
@@
    Taxonomy. We categorized the risks LLMs face based on their life cycle. Other surveys lacked a fine-grained taxonomy, making it difficult to distinguish the characteristics of various risks.
@@
    Unique privacy/security. Regarding privacy/security risks, we focused on the unique risks to LLMs. Additionally, we explored the common risks associated with all language models. In response to these risks, we summarized potential countermeasures. However, most surveys lacked depth and comprehensiveness in this analysis, listing attacks and defense methods.
@@
    Other unique privacy/security scenarios. We incorporated LLMs within three additional scenarios: federated learning, machine unlearning, and watermarking. Since these scenarios involve privacy/security threats, we explored them in Section 9 and provided a systematic study that most surveys overlooked.
@@ 1.3. Contributions of this survey
LLMs possess strong emergent abilities, which are applied across many industries. However, numerous vulnerabilities in their life cycle pose significant privacy and security risks. These risks seriously hinder the application and expansion of LLMs. Hence, we aim to analyze, categorize, and summarize these privacy and security issues. Specifically, we propose a novel taxonomy for these risks, providing a clear and comprehensive analysis of their goals, causes, and implementation methods. We hope this survey provides researchers with feasible research directions and challenges. The main contributions are as follows.
@@
    Taking the LLMs’ life cycle as a clue, we considered risks and countermeasures in five different scenarios from those of traditional models. These scenarios include pre-training, fine-tuning, RAG systems, deployment and LLM-based agents.
@@
    For each scenario, we highlighted the differences in privacy and security risks between LLMs and traditional models. Specifically, we described unique risks to LLMs and common risks to all models. For each risk, we detailed its attack goal and capacity, and reviewed related research using the attack method. Additionally, we analyzed the mapping between risks and countermeasures, listing potential defense methods.
@@ 2.1. Definition of LLM
LLMs represent a revolution in the field of NLP. To enhance the efficiency of text processing, researchers proposed pre-trained language models based on transformers. Google released the BERT model, which uses bidirectional transformers, solving downstream tasks in the ’pre-train + fine-tune’ paradigm. Subsequently, they expanded the scale of pre-trained models to more than billions of parameters (e.g., GPT-3) and introduced novel techniques. Large-scale pre-trained models showcase remarkable emergent abilities not found in regular-scale models, capable of handling unseen tasks through in-context learning (Xu et al., 2024a) (i.e., without retraining) and instruction tuning (Shu et al., 2023) (i.e., lightweight fine-tuning).
Recent studies have summarized four key characteristics that LLMs should possess (Zhao et al., 2023c; Xu et al., 2024b). First, an LLM should understand various texts, such as training data and natural language instructions. Second, it should solve unseen NLP tasks without updating the model parameters. Third, it should generate high-quality texts that align with human expectations when provided with designed prompts. Fourth, an LLM should demonstrate contextual awareness, considering specific factors such as domain expertise. Moreover, Wei et al. (Wei et al., 2022) found that language models with more than 1 billion parameters exhibit significant performance improvements on multiple NLP tasks. Therefore, we believe that LLMs must possess more than a billion parameters. Numerous institutions have developed LLMs with these characteristics.
@@ 2.2. Traditional privacy and security risks
Recent research on privacy and security risks in artificial intelligence has primarily focused on small-scale models.
Regarding privacy risks, the life cycle of small-scale models contains confidential information such as raw data and model details. Leakage of this information could lead to severe economic losses (Ye et al., 2024). Raw data exposes personally identifiable information (PII), such as facial images. Reconstruction attacks (Morris et al., 2023b) and model inversion attacks (Ye et al., 2024) can extract raw data using gradients or logits. Additionally, membership and attribute information are sensitive. For example, in medical tasks, adversaries can use membership inference attacks (Ye et al., 2022) to determine if an input belongs to the training set, revealing some users’ health conditions. Model details have significant commercial value and are vulnerable to model extraction attacks (Li et al., 2023e), which target black-box victim models to obtain substitute counterparts or partial model information via multiple queries. Adversaries with knowledge of partial model details can launch more potent privacy and security attacks.
Regarding security risks, small-scale models face poisoning attacks (Wan et al., 2023), which compromise model utility by modifying the training data. A backdoor attack is a variant of poisoning attacks (Wang et al., 2023a; Ma et al., 2023). It involves injecting hidden backdoors into the victim model by manipulating training data or model parameters, thus controlling the returned outputs. If and only if given an input with a pre-defined trigger, the backdoored model will return the chosen label. During inference, adversarial example attacks (Guo et al., 2021) craft adversarial inputs by adding imperceptible perturbations, causing incorrect predictions. In summary, these security attacks can compromise model utility and integrity, severely threatening public safety in practical applications.
The life cycle of LLMs shares similarities with, yet also differs from, that of small-scale models. Building on traditional privacy and security risks, we outline various potential risks LLMs face in different scenarios, as illustrated in Figure 1. Each scenario involves unique data types and implementation processes, presenting different risks and their corresponding countermeasures. We will begin by discussing each scenario and its associated threat models.
Figure 1. The pipeline of the survey. For each threat scenario, the first column lists the data type used, and the second column describes the process applied. The text boxes indicate unique data types and processes to LLMs. The fourth and fifth columns detail the extended risks and countermeasures, respectively. Underlined texts represent unique risks for LLMs, while other texts indicate common risks for all language models.
@@ 3. Threat Scenarios for LLMs
Although many institutions have disclosed the implementation methods of their LLMs, some details remain unknown. We have reviewed the literature and divided the life cycle of an LLM into five scenarios with finer granularity rather than just the training and inference phases. Figure 1 illustrates these unique threat scenarios: pre-training LLMs, fine-tuning LLMs, RAG systems, deploying LLMs, and deploying LLM-based agents. We then list extended risks for each scenario, using underlined texts to highlight unique risks for LLMs.
@@ 3.1. Pre-training LLMs
In this scenario, model developers collect a large corpus as a pre-training dataset, including books (Zhu et al., 2015), web pages (e.g., Wikipedia), conversational texts (e.g., Reddit), and code (e.g., Stack Exchange). They then use large-scale, Transformer-based networks and advanced training algorithms, enabling the models to learn rich language knowledge from vast amounts of unlabeled texts. After obtaining the pre-trained LLM, developers upload it to open-source community platforms to gain profits, as shown in Figure 2. In this context, we consider three malicious entities: contributors, developers, and users.
@@
    Malicious contributors. Unlike small-scale models, the corpora involved in pre-training LLMs are so large that developers cannot audit all the data, resulting in the inevitable inclusion of negative texts (e.g., toxic data and private data). These negative texts directly impact the safety of LLMs. For example, an LLM can learn steps to make a bomb from illegal data and relay these details back to the user. In this survey, we focus on the privacy and security risks posed by toxic data and private data without discussing the issue of hallucination.
@@
    Malicious developers. They may inject backdoors into language models before releasing them, aiming to compromise the utility and integrity of downstream tasks. If victim users download and deploy a compromised model, an attacker who knows the trigger can easily activate the hidden backdoor, thus manipulating the compromised model.
@@
    Malicious users. After downloading public models, they gain access to the model’s information except for the training data, effectively becoming white-box attackers. Consequently, these users can perform inference and data extraction attacks in a white-box setting.
@@ 3.2. Fine-tuning LLMs
In this scenario, users customize LLMs for specific NLP tasks. They download pre-trained LLMs from open-source platforms and fine-tune them on customized datasets. There are three fine-tuning methods: supervised learning, instruction-tuning, and alignment-tuning. The first method is the commonly used training algorithm. For the second method, the instruction is in natural language format, containing a task description, an optional demonstration, and an input-output pair (Wang et al., 2023b). Through a sequence-to-sequence loss, instruction-tuning helps LLMs understand and generalize to unseen tasks. The third method aligns LLMs’ outputs with human preferences, such as usefulness, honesty, and harmlessness. To meet these goals, Ziegler et al. (Ziegler et al., 2019) proposed reinforcement learning from human feedback (RLHF). Figure 5 illustrates two types of malicious entities: contributors and third parties.
@@
    Malicious contributors. Generally, users need to collect specific samples used to fine-tune downstream tasks. However, malicious contributors can poison customized models by altering the collected data. In this scenario, the adversary can modify a fraction of the contributed data without access to the fine-tuning process.
@@
    Malicious third-parties. When outsourcing customized LLMs, users share their local data with third-party trainers who possess computational resources and expertise. However, malicious trainers can poison these customized LLMs to manipulate their responses before delivering them to users. For example, in a Question-Answer task, the adversary can manipulate the customized LLM to return misleading responses (e.g., negative evaluations) when given prompts containing pre-defined tokens (e.g., celebrity names). Compared to small-scale models, fine-tuning LLMs involves two unique techniques: instruction-tuning and alignment-tuning. Therefore, malicious trainers pose two unique risks to LLMs: poisoning instruction-tuning and RLHF. Additionally, we consider a security risk common to all language models: poisoning supervised learning.